The difference between the triple of a number and a quarter of another number is 13 and the other is double the first number plus twice the second is 18. Find the numbers
Answer by Theo(13342)

your two equations are:

3x - y/4 = 13

2x + 2y = 18

solve for y in the first equation to get y = 12x - 52.

replace y with 12x - 52 in the second equation to get:

2x + 2 * (12x - 52) = 18

solve for x to get x = 122/26

solve for y in the first equation to get y = 112/26.

your solution is that x = 122/26 and y = 112/26.

replace x and y with these values in both the original equations and you will find that they are true.

the mathematics is cumbersome.

if you have a calculator with more than one memory, it pays to do the math using your calculator and storing the results.

doing it that way, i got:

x = 4.692307692

y = 4.307692308